OpenStudy (anonymous):

if a circle is tangent to a line through (-2,5) and (4,3) find the slope of the radius to the point of contact

OpenStudy (anonymous):

@ganeshie8

OpenStudy (bradely):

tangent of the circle and slope of the radius to the point of contact are perpendicular Slope of the tangent , m = (3-5)/(4-(-2)) = -2/6 =-1/3 Slope of the radius to the point of contact is -1/m = 3 Source: mathskey.com